OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

OpenStudy (anonymous):

Let y=0, 0=-2x-3, 3=-2x, x=-1.5 (x intercept) - (-1.5 or -3/2, 0) Let x=0, y=-2(0)-3, y=0-3, y=-3 (y intercept) - (0,-3)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem